# Clinical Data Goes MEDS? Let’s OWL make sense of it

### v1

@@ 4 Discussion
Our testing on the neurovasc dataset shows that MEDS-OWL can capture the MEDS event and their provenance, including their conversion to RDF itself, while remaining compact and pragmatic. The RDF graph produced is SHACL-validated, supporting semantic and structural consistency. Moreover, exporting MEDS datasets as KGs enables the direct application of graph representation learning techniques, thereby opening new opportunities for downstream applications such as graph-based ML.
@@
Compared to existing approaches for representing clinical data in RDF, such as FHIR-RDF, SPHN, and CARE-SM, MEDS-OWL addresses several gaps. Unlike FHIR-RDF, which suffers from loosely specified semantics, or SPHN and CARE-SM, which are tightly coupled to specific initiatives or specific aims, MEDS-OWL follows the minimality of MEDS, and its ML-oriented design, plus a well defined semantics with OWL and reused ontologies. This facilitates interoperability across heterogeneous datasets through explicit mapping points (e.g., meds:parentCode or skos:exactMatch).
@@
We acknowledge several limitations of the current study. First, the preliminary testing we achieved is limited to the transformation of a single dataset. Transforming other datasets to MEDS-OWL, computing associated stats, and conducting an error analysis of SHACL inconsistencies potentially arising would enable to validate further the generalizability of our approach. Indeed, we did evaluated the reasoning performance or scalability of the generated KGs; processing large, real-world datasets may reveal bottlenecks in RDF generation, triple-store ingestion, or reasoning that will require optimization. Moreover, the current MEDS-OWL ontology does not capture certain operational constraints of MEDS datasets. Specifically, MEDS requires subject contiguity, what means that all the features associated to a single subject should appear in a single sort and should be temporally sorted. While temporal ordering becomes largely irrelevant when converting tabular MEDS data to RDF graphs, subject contiguity can be partially represented by associating each subject with a specific ML partition via the meds:SubjectSplit class. Importantly, a single subject may participate in multiple independent splits (e.g., for repeated experiments or cross-validation); however, within each split, the subject can belong to only one subset (e.g., training, validation, or test). This ensures that partitioning is mutually exclusive at the level of each experiment while still supporting multiple, independent splits for variability analysis. To address this limitation, future releases of MEDS-OWL will incorporate enhanced axioms, potentially using disjointness rules or qualified cardinality restrictions in description logics, thereby allowing the ontology to more accurately represent MEDS operational constraints.
@@
To substantiate the claimed benefits and broaden the applicability of the ontology, we established an evaluation agenda and identified four steps. (1) Quantifying round-trip fidelity to measure information preservation and loss due to transformations. (2) Aggregating SHACL validation statistics across multiple, heterogeneous datasets to characterize common data-quality issues and to demonstrate how schema validation reduces downstream errors. (3) Profile scalability and performance on large, real-world datasets. Specifically, we plan to apply our framework to the MIMIC-IV dataset [13]. Compared to the synthetic neurovasc dataset, MIMIC-IV is substantially larger and more complex, providing an opportunity to assess the semantic fidelity and the scalability of the graph generation process. (4) Compare predictive analysis on MEDS-OWL data with graph-based ML models, such as GCNs.
@@
Planned extensions of the MEDS-OWL ontology and associated tooling include greater reuse of external ontologies to replace remaining MEDS-specific constructs, thereby reducing bespoke terms and enhancing interoperability. We aim to integrate temporal ontologies (e.g., OWL-Time) and pattern-based representations to more accurately model intervals and episode boundaries, as well as adopt unit ontologies to standardize numeric values and support unit conversion and aggregation. Hierarchical relations among codes, currently captured via meds:parentCode (see Table 3), will be formalized using skos:broader to facilitate cross-vocabulary linking, with optional use of rdfs:subClassOf for implementations that model codes as OWL classes. These enhancements aim to increase semantic richness and usability of MEDS-OWL for KG, enabling systematic cross-vocabulary linking.
@@ 5 Conclusion
We have presented MEDS-OWL, an ontology that formalizes the MEDS event model using Semantic Web standards, along with the meds2rdf library for converting MEDS datasets into RDF graphs compliant with MEDS-OWL. Our framework preserves MEDS’s ML–oriented representation of medical event sequences while enhancing it with a rich semantic structure and explicit links to biomedical and provenance ontologies. This enables symbolic reasoning, interoperability, and integration with other knowledge-based systems.
By providing a reliable transformation from MEDS to RDF, our framework lays the groundwork for downstream prediction tasks using graph-based ML models. Preliminary experiments on the synthetic neurovasc dataset demonstrate that MEDS-OWL and meds2rdf can accurately transform MEDS events and their associated metadata. Released under an open-source license, these tools promote FAIR-aligned sharing and reuse of medical event data.
In future work, we plan to investigate how RDF-based representations affect predictive performance and to further explore the potential of graph representation learning approaches.
@@ Acknowledgments
This work is supported by CombO (Health Data Hub) project and the the Agence Nationale de la Recherche under the France 2030 program, reference ANR-22-PESN-0007 ShareFAIR, and ANR-22-PESN-0008 NEUROVASC.

### v2

@@ 4 Discussion
Our evaluation on two complementary datasets, the synthetic neurovasc cohort and a real-world subset of MIMIC-IV, demonstrates that MEDS-OWL can faithfully represent MEDS event data and associated provenance in RDF while remaining compact and operationally pragmatic. The resulting RDF graphs are SHACL-validated, supporting semantic and structural consistency. Moreover, exporting MEDS datasets as KGs enables the direct application of graph representation learning techniques, thereby opening new opportunities for downstream applications such as graph-based ML. At the same time, the KG representation allows us to leverage the strengths of both paradigms: the event-centric, ML-oriented design of MEDS and the semantic expressivity of KGs. By explicitly modeling entities, relations, and provenance, MEDS-OWL provides semantic enrichment of the underlying datasets, making complex, multi-relational dependencies between clinical events accessible to downstream prediction models. Importantly, because MEDS-OWL preserves the temporal structure and ordering of events, it naturally supports the construction of Temporal Knowledge Graphs (TKGs), enabling longitudinal modeling of patient trajectories and opening new directions for temporally aware clinical prediction tasks.
@@
Compared to existing approaches for representing clinical data in RDF, such as FHIR-RDF, SPHN, and CARE-SM, MEDS-OWL addresses several gaps. Unlike FHIR-RDF, which suffers from loosely specified semantics, or SPHN and CARE-SM, which are tightly coupled to specific initiatives or aims, MEDS-OWL follows the minimality of MEDS, its ML-oriented design, and a well-defined OWL semantics reusing ontologies. This facilitates interoperability across heterogeneous datasets through explicit mapping points (e.g., meds:parentCode or skos:exactMatch).
@@
The dual evaluation highlights both strengths and current limitations. While the neurovasc dataset illustrates a test on a controlled synthetic cohort, the MIMIC-IV conversion demonstrates that the approach scales to large, heterogeneous, real-world data. The latter confirms that meds2rdf can process high event volumes while preserving semantic consistency. It also emphasizes the need for careful performance profiling and analysis to avoid SHACL inconsistencies when handling millions of triples and thousands of distinct clinical codes. Currently, we have not conducted an in-depth analysis of reasoning performance or triple-store ingestion costs, which remain important directions for future work in large-scale research applications. Moreover, the current MEDS-OWL ontology does not capture certain operational constraints of MEDS datasets. Specifically, MEDS requires subject contiguity, which means that all the features associated with a single subject should appear in a single sort and should be temporally sorted. While temporal ordering becomes largely irrelevant when converting tabular MEDS data to RDF graphs, subject contiguity can be partially represented by associating each subject with a specific ML partition via the meds:SubjectSplit class. Importantly, a single subject may participate in multiple independent splits (e.g., for repeated experiments or cross-validation); however, within each split, the subject can belong to only one subset (e.g., training, validation, or test). This ensures that partitioning is mutually exclusive at the level of each experiment while still supporting multiple, independent splits for variability analysis. To address this limitation, future releases of MEDS-OWL will incorporate enhanced axioms, potentially using disjointness rules or qualified cardinality restrictions in description logics, thereby allowing the ontology to more accurately represent MEDS operational constraints.
@@
To substantiate the claimed benefits and broaden the applicability of the ontology, we established an evaluation agenda and identified four steps. (1) Quantifying round-trip fidelity to measure information preservation and loss due to transformations. (2) Aggregating SHACL validation statistics across multiple, heterogeneous datasets to characterize common data-quality issues and to demonstrate how schema validation reduces downstream errors. (3) Profiling performance on large MEDS corpora will facilitate the identification of bottlenecks and enable a rigorous assessment of semantic fidelity and graph generation scalability. (4) Compare predictive analysis on MEDS-OWL data with graph-based ML models, such as GCNs.
@@
Planned extensions of the MEDS-OWL ontology and associated tooling include greater reuse of external ontologies to replace remaining MEDS-specific constructs, thereby reducing bespoke terms and enhancing interoperability. We aim to integrate temporal ontologies (e.g., OWL-Time) and pattern-based representations to more accurately model intervals and episode boundaries, while also adopting unit ontologies to standardize numeric values and support unit conversion and aggregation. Hierarchical relations among codes, currently captured via meds:parentCode (see Table 4), will be formalized using skos:broader to facilitate cross-vocabulary linking, with optional use of rdfs:subClassOf for implementations that model codes as OWL classes. These enhancements aim to increase the semantic richness and usability of MEDS-OWL for KG, enabling systematic cross-vocabulary linking.
@@ 5 Conclusion
We have presented MEDS-OWL, an ontology that formalizes the MEDS event model using Semantic Web standards, along with the meds2rdf library for converting MEDS datasets into RDF graphs compliant with MEDS-OWL. Our framework preserves MEDS’s ML–oriented representation of medical event sequences while enhancing it with a rich semantic structure and explicit links to biomedical and provenance ontologies. This combination enables symbolic reasoning, interoperability across heterogeneous data sources, and seamless integration with knowledge-based systems.
Preliminary experiments on the synthetic neurovasc dataset and a real-world subset of MIMIC-IV demonstrate that MEDS-OWL and meds2rdf can accurately transform MEDS events and their associated metadata into semantically validated RDF graphs. Released under an open-source license, these tools promote FAIR-aligned sharing and reuse of medical event data.
By providing a reliable transformation from MEDS to RDF, our framework lays the groundwork for downstream prediction tasks using graph-based ML models. Representing MEDS datasets as KGs enables semantic enrichment of clinical events and their use across multiple analytical tasks. Moreover, because MEDS-OWL preserves temporal ordering, it naturally supports the construction of TKGs, opening new avenues for longitudinal patient modeling and temporally aware prediction tasks that go beyond conventional tabular approaches.
In future work, we plan to systematically compare graph-based ML models with classical ML approaches to assess whether RDF-based, relational representations provide measurable improvements in predictive performance over traditional tabular methods. Additionally, we aim to evaluate the round-trip fidelity of the framework on other public biomedical datasets formatted in MEDS, thereby assessing information preservation across transformations. Finally, we plan to further align MEDS-OWL with widely adopted biomedical ontologies to strengthen semantic interoperability and facilitate broader reuse across biomedical KGs ecosystems.
Reproducibility Statement
All experiments and dataset conversions presented in the accompanying paper are fully reproducible using the workflows provided in this repository: https://github.com/albertomarfoglia/meds-to-owl-examples. The end-to-end pipelines for the neurovasc synthetic dataset and the MIMIC-IV Demo MEDS dataset are implemented as executable, notebook-based workflows (main.ipynb). They rely only on open-source tools and publicly available data or scripts. The repository specifies all software dependencies and configuration files. Moreover, randomized steps, such as synthetic data generation or sampling, are executed with fixed seeds to ensure repeatability.
@@ Acknowledgments
This work is supported by CombO (Health Data Hub) project and the Agence Nationale de la Recherche under the France 2030 program, reference ANR-22-PESN-0007 ShareFAIR, and ANR-22-PESN-0008 NEUROVASC.